# Learning to Generate Unit Tests for Automated Debugging

### v1

@@ 3.3 UTDebug: Debugging with Generated Unit Tests
A crucial difference between using generated unit tests (as opposed to human-generated or gold unit tests) is the degree of noise in unit test feedback. Despite training models – via UTGen or otherwise – a generated unit test may not be 100% accurate. This manifests in two ways: 1) the generated unit test input is not failing for a given code under debugging , i.e., ; 2) the generated unit test output is inaccurate, i.e., not consistent with what a gold solution to the task description would yield (). Both these types of errors can negatively impact downstream utility of unit tests (in our context, ability to debug) in different ways as illustrated in Figure 3 (left). First, a lack of adversarial input could result in a faulty code (albeit with subtle errors) being misclassified as correct, and thus, erroneously taken out of the debugging process. Next, incorrect unit test outputs could also lead to false positive, with subsequent edits introducing errors to an otherwise correct code. Moreover, even if the candidate code contains bugs, an incorrect unit test output can result in incorrect feedback given to the model during debugging, which in turn could be wasteful or detrimental to downstream performance. These considerations motivate our RQ3 on how best to incorporate potentially noisy feedback. Below we propose UTDebug, which contains two effective ways of mitigating noisy feedback from automatically generated unit tests (with additional details in Appendix C):
@@
Boosting Output Accuracy via Test-Time Scaling. Building on past work that has shown the benefits of scaling up inference-time compute (Wang et al., 2022; Lightman et al., 2023; Snell et al., 2024) for LLMs, we first improve output accuracy by allocating additional computation to the problem. Specifically, we use self-consistency (SC; Wang et al., 2022), whereby, for a given UT input, we sample output completions (including CoT rationales) and take the most common final UT output (majority vote) as the final answer, as shown in Figure 3 (top-right). To further boost output accuracy, we upsample UT inputs and only retain those where the final answer gets over 50% of the votes (i.e., 4 votes), discarding the unit test otherwise. Our results in Section 5.3 show that this boosts accuracy by up to 11.4%.
@@
Back-Tracking and Cross-Validation. In addition to enabling test-time scaling to improve output accuracy, in settings with noisy feedback, it is important to know when to abstain or backtrack, e.g., discarding edits made to a candidate code. This can be beneficial when the generated unit test has an incorrect output, and therefore, the feedback is incorrect. It can also come into play when the feedback is sound but is not correctly incorporated by the model. Inspired by test-driven development with human programmers debugging a failing unit test, we accept changes only if the revised code passes the previously failing UT. Moreover, in practice, developers typically use a suite of multiple UTs; when changes are made to a codebase, all UTs are run and errors are logged. Such a suite can inform developers when a change or fix in one part of the code has caused errors in another part. In the context of debugging, this process can help prevent overfitting, whereby the debugging agent modifies code to make a single unit test pass, but thereby causes other tests to fail, leading to one unit test passing, but making the code unusable generally. Therefore, when debugging via UTGen, in each round we generate UTs. We use one to provide feedback to the debugging LLM, and accept the debugger’s edits only if the pass rate on the entire set of unit tests improves; if it does not, we backtrack. This is illustrated in Figure 3 (b) as the edits in round 1 are discarded and backtracked whereas the edits in round 2 improve the entire test suite and are accepted.
@@ 4 Experimental Setup
We use debugging datasets based on popular LLM coding benchmarks, HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) along with their extended evaluation suites with more unit tests as proposed by (Liu et al., 2024). We describe their construction below, with further details in Appendix A.
HE+Fix. We start with the HumanEvalFix dataset containing human-introduced errors to gold solutions proposed by Muennighoff et al. (2024). However, this dataset only uses minimal unit tests to evaluate code correctness (from HumanEval) which has shown to be unreliable (Li et al., 2023), as it can miss errors due to low coverage. Therefore, we filter for problems that overlap with EvalPlus (Liu et al., 2024), which contains over 80 more unit tests. This yields 158 problems that each have a faulty solution and an expanded suite of private UTs – i.e., UTs not used for debugging – for evaluation.
@@
MBPP+Fix. We construct a debugging dataset based on the MBPP+ benchmark (Austin et al., 2021; Liu et al., 2024; Ni et al., 2024). To obtain faulty codes, we sample 16 solutions for each problem from different 7-8B scale models and filter for incorrect solutions based on the private unit tests (described in Appendix A). Then, we select one faulty solution at random corresponding to each problem. This results in a total of 325 problems with incorrect solutions representing realistic errors made by LLM coders. We find such errors are generally more challenging for models to debug (cf. Section 5).

### v2

@@ 3.3 Generated Unit Tests for Debugging
A key difference when using generated unit tests, as opposed to human-generated UTs, is the degree of noise in the feedback. Despite training models via UTGen or other methods, a generated unit test may not be 100% accurate. This can manifest in two ways: 1) the generated UT input does not fail for the code under debugging , i.e., ; 2) the generated UT output is inaccurate, i.e., not consistent with what a gold solution would yield (). Both types of errors can negatively impact the utility of unit tests for debugging, as shown in Fig. 3 (left). A non-adversarial input might result in faulty code being misclassified as correct and prematurely removed from debugging. Additionally, incorrect outputs can cause false positives, introducing errors to otherwise correct code. Even if the candidate code has bugs, incorrect UT outputs can lead to incorrect feedback, thereby, degrading performance. These issues motivate our RQ3 to incorporate noisy feedback. We propose UTDebug, which includes two ways to mitigate noisy feedback from automatically generated unit tests (detailed in Appendix C):
Boosting Output Accuracy via Test-Time Scaling.
@@
Building on past work that has shown the benefits of scaling up inference-time compute for LLMs (Wang et al., 2022; Lightman et al., 2023; Snell et al., 2024), we improve UT output accuracy by allocating additional computation to the problem. Specifically, we use self-consistency (SC; Wang et al., 2022), whereby, for a given UT input, we sample output completions (including CoT rationales) and take the most common final UT output (majority vote) as the final answer, as shown in Fig. 3 (top-right). To further boost output accuracy, consistent with Prasad et al. (2024), we upsample UT inputs and only retain those where the final answer gets over 50% of the votes (i.e., 4 votes), discarding the unit test otherwise.
Back-Tracking and Validation.
@@
To handle noisy feedback, it is crucial to know when to abstain or backtrack, e.g., discard edits. This is useful when generated UT outputs are incorrect, leading to faulty feedback, or when correct feedback is not incorporated by the model. Inspired by test-driven development, we accept changes only if the revised code passes the previously failing UT. Moreover, developers often use multiple UTs to detect errors from changes in other parts of the code. This helps prevent overfitting, where code is modified to pass a single unit test but fails others. Therefore, in each debugging round via UTGen, we generate UTs. We use one for debugging feedback and accept edits only if the pass rate on the entire set improves (validation); otherwise, we backtrack. This process is shown in Fig. 3 (b), where edits in round 1 are discarded, but those in round 2 improve the test suite as a whole and are accepted.
@@ 4 Experimental Setup
We use debugging datasets based on popular LLM coding benchmarks, HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) along with their extended evaluation suites with more unit tests as proposed by Liu et al. (2024b). We describe their construction below, with further details in Sec. A.1.
HE+Fix. We take HumanEvalFix (Muennighoff et al., 2024), which has human-introduced errors, and following Ni et al. (2024), augment each problem’s test suite with overlapping problems from EvalPlus Liu et al. (2024b). This yields 158 problems, each with a faulty solution and an expanded suite of private UTs for evaluation – i.e., UTs not used for debugging.
@@
MBPP+Fix. We construct a debugging dataset based on the MBPP+ benchmark Austin et al. (2021); Liu et al. (2024b); Ni et al. (2024). To obtain faulty codes, we sample 16 solutions for each problem from different 7-8B scale models and filter for incorrect solutions based on the private unit tests (described in Sec. A.1). Then, we sample one faulty solution for each problem, resulting in 325 problems with incorrect solutions, representing realistic errors made by LLMs. We find such errors are generally harder for models to debug (cf. Sec. 5).
MBPP+Fix (Hard). Additionally, we create a different split from MBPP+ with more subtle errors, logical flaws and missing corner cases making it harder to debug. To this end, we follow a similar generation and filtering process as above but only retain faulty code that passes between of unit tests. This results in a total of 170 problems with faulty codes.
Evaluation Metrics.
Below we describe three intrinsic evaluation metrics for unit test generation (with additional details in Sec. A.2):
Attack Rate. We measure the attacking ability of a UT generator by matching how often the output of a gold reference solution for input differs from that of the buggy code .
Output Accuracy. This metric measures how often the output of a generated unit test is consistent with the problem description, i.e., generates the same output as the reference gold code .
Accuracy Attack. This metric combines both attack rate and output accuracy and represents how often a unit test generator generates a useful (i.e., failing) unit test for a given target code while also predicting the output correctly.
We rely on intrinsic metrics during development on a validation split from MBPP+ data, e.g., for checkpoint selection and prompt design. Then, we measure the utility and effectiveness of UTGen by computing the pass@1 code accuracy, i.e., the percentage of codes passing all unit tests on code debugging and generation tasks.
Baselines.
We compare UTGen against the following (prompts in Appendix E):
No UT feedback: In this baseline, we use self-generated or self-critique feedback, following prior work Madaan et al. (2024); Shinn et al. (2024); Chen et al. (2023b). Specifically, we prompt the model to generate an explanation about the bugs in the target code in addition to determining if the target code is correct, i.e., does not have any more bugs.
Prompted Failing UTs: Using the same prompts as UTGen without any training, we prompt an LLM to generate a failing UT (with a correct output) given the description and the target code. Here, we sample .
5 Results and Analysis
5.1 Intrinsic Evaluation of UT Generation
In order to study the inherent UT generation abilities of different models and baselines, we use the intrinsic metrics defined by RQ1 and outlined in Sec. 4 in Sec. 5.1 on the most challenging debugging task MBPP+Fix (Hard) averaged over 3 runs.
Trade-offs in Attack Rate and Output Accuracy.
In Sec. 5.1, we benchmark the zero-shot abilities of different models on UT generation, as well as our improved training method (corresponding to RQ2). Here, we observe a clear tradeoff: prompted and random unit test generation each optimizing output accuracy or attack rate at the cost of the other metric. While randomly sampled UTs have relatively higher output accuracy (i.e., the output of the unit test is correct according to the problem description), the random baseline often lacks the ability to generate UTs with inputs that trigger errors (failing UTs). This can be explained by the fact that it is not conditioned on the faulty code . On the other hand, when models are prompted to generate a UT that breaks the faulty code in the prompted UT baseline, they generally achieve a higher attack rate; however, they lag in terms of output accuracy. For instance, in the case of Qwen2.5 7B, switching from random sampling to prompting failing UTs improves the attack rate by 13.53% but decreases the output accuracy by 16.47%, hence leading to comparable “Acc. Attack” scores (13.53% vs 14.71%). Taken together, this suggests the presence of competing desirable traits of UT generators (cf. Sec. 3.1; RQ1): without training, one baseline often generates trivial unit tests for which the output can easily be predicted (high output accuracy, low attack rate), while the other generates challenging unit tests for which it cannot produce the expected answer (low output accuracy, high attack rate). It also indicates that while models can generate challenging unit test inputs that result in code failure, models struggle with reasoning over correct outputs of failing UT inputs.
l c c c c
Model Method Attack Rate Out Acc. Acc. Attack
Llama3
8B Random 30.98 46.86 10.24
Prompted 38.04 37.10 11.51
UTGen 41.18 48.24 16.57
Llama3.1
8B Random 21.76 39.80 9.02
Prompted 40.59 26.86 9.61
UTGen 41.37 47.75 16.67
Qwen2.5
7B Random 26.27 57.25 13.53
Prompted 39.80 40.78 14.71
UTGen 41.96 58.04 20.20
Qwen2.5
32B Random 25.29 47.25 11.37
Prompted 51.96 48.43 29.41
UTGen 56.08 59.22 34.71
UTGen is Most Effective at Unit Test Generation.
We propose to break through this trade-off (thereby addressing RQ2, “How can we improve LLMs’ UT generation abilities?”) by training models with UTGen, where we directly supervise models to have both high output accuracy and high attack rate. Ideally, UTGen should lead to models that generate unit tests in the “sweet spot” of difficulty, where they are hard enough to trigger errors but not so hard that their output cannot be predicted. Sec. 5.1 shows that models trained with UTGen consistently rank highest at jointly generating failing UT inputs with correct outputs (as measured via “Acc. Attack” in Sec. 5.1). For instance, on the strongest coding model in our setting, Qwen2.5 32B, UTGen obtains the highest attack rate, and output accuracy as well as improves Acc. Attack score by 5.3% (absolute) over failing UTs generated by prompting and by 23.34% over randomly generated UTs. Similarly, Llama3.1 improves (in terms of attack rate) over randomly-sampled UTs by 19.61% and the output accuracy of prompted failing UTs by 20.89%, ultimately improving the joint “Acc. Attack” score by up to 7.65%. We report the intrinsic metrics for HE+Fix and MBPP+Fix in Appendix D.
| Model | UT Method | HE+Fix | MBPP+Fix | (Hard) |
| Llama3 8B | No UT | 27.22 | 16.31 | 11.76 |
| Random | 51.90 | 30.46 | 17.06 | |
| Prompted | 51.90 | 28.92 | 22.94 | |
| UTGen | 53.80 | 37.54 | 28.82 | |
| Llama3.1 8B | No UT | 31.65 | 10.15 | 11.18 |
| Random | 62.03 | 33.54 | 13.53 | |
| Prompted | 56.33 | 28.00 | 24.71 | |
| UTGen | 67.09 | 36.92 | 28.23 | |
| Qwen2.5 7B | No UT | 52.53 | 23.08 | 16.47 |
| Random | 79.75 | 34.77 | 17.06 | |
| Prompted | 75.32 | 32.92 | 24.12 | |
| UTGen | 82.91 | 37.54 | 29.41 | |
| Qwen2.5 32B | No UT | 79.11 | 39.08 | 32.94 |
| Random | 84.81 | 49.54 | 22.94 | |
| Prompted | 85.44 | 50.77 | 40.59 | |
| UTGen | 88.61 | 54.15 | 45.29 |
5.2 Generated UTs for Code Debugging
UTs from UTGen are Best for UTDebug.
Addressing RQ3, we measure how effective each type of UT generator is in a downstream debugging evaluation. From Sec. 5.1, we observe that across different LLM families, multi-turn debugging with UTDebug is more effective when using generated UTs from UTGen than debugging with UTs generated by the baselines, as well as debugging without feedback. For instance, with Qwen2.5 32B, after 3 rounds of debugging on MBPP+Fix, UTGen improves over debugging with randomly-sampled UTs by 4.61%, debugging by failing UTs by 3.38%, and debugging without UT feedback by 15.07%. Moreover, on the more challenging MBPP+Fix (Hard) split, UTGen improves over randomly-sampled UTs by 22.35% and the baseline without UT feedback by 12.35%. Given that we use the same underlying LLM and similar feedback templates, the results in Sec. 5.1 show that UTs generated by UTGen provide the most useful and effective feedback. Lastly, we find that LLM debugging without any UT feedback is least effective across model families, thus, establishing that despite noise and issues discussed in Sec. 3.3 and Fig. 3, UTDebug with UT-based feedback is more effective than self-generated feedback. This aligns with past work indicating that LLMs are poor at self-critique Huang et al. (2024); Stechly et al. (2024); Chen et al. (2024).
Debugging Difficulty Varies Across Datasets.
Comparing the post-debugging accuracy across the three datasets in Sec. 5.1 suggests different difficulty levels when it comes to LLM debugging. The relatively high accuracies on HE+Fix (especially with Qwen2.5) suggest that human-introduced errors in that dataset are obvious and thus too easy for the models to discover and debug. This is also corroborated by the fact that starting codes in this dataset have a pass rate of 0%, i.e., fail on all unit tests, suggesting high-level coarse-grained errors, e.g., syntax errors that prevent execution, or incorrect function references. On the other hand, MBPP+Fix (Hard) (with an initial pass rate of 75%) appears to be the hardest to debug, with the lowest overall post-debugging accuracy across models – up to a difference of 53.5% in final accuracies of HumanEvalFix and MBPP+Fix (Hard) for UTGen based on Qwen2.5 7B. This in turn suggests that LLMs still struggle with identifying and fixing subtle flaws in generated code, especially in scenarios involving corner cases wherein the model passes some unit tests and fails at the rest.
5.3 Generated UTs for Code Generation
| Model | Judge Method | HE+ | MBPP+ |
| Llama3 8B | -† | 54.43 | 56.00 |
| RM Liu et al. (2024a) | 60.76 | 59.47 | |
| Random UT | 61.39 | 58.93 | |
| Prompted UT | 63.29 | 60.00 | |
| UTGen | 65.12 | 61.60 | |
| Qwen2.5 7B | -† | 81.64 | 65.06 |
| RM Liu et al. (2024a) | 83.54 | 66.67 | |
| Random UT | 85.44 | 68.80 | |
| Prompted UT | 85.71 | 67.47 | |
| UTGen | 87.97 | 70.13 |
Setup.
We further demonstrate the effectiveness of generated UTs at judging code correctness via Best-of- sampling (). Given a problem , we sample different code solutions, and for each solution, we generate 3 UTs (via UTGen and baselines in Sec. 4 with test-time scaling). In the end, we collate all generated UTs from each method (removing any duplicates) and choose the generated code solution that passes the most generated UTs. In other words, we use UTGen to rerank 10 generated solutions. In addition to comparing to the random and prompt UT generation baselines, we also rerank according to a state-of-the-art 8B scale reward model Liu et al. (2024a) on HE+ and MBPP+ (two standard code-generation datasets).
Results.
Based on results in Table 3, we find that:
Generated UTs Outperform External RMs. In all four settings (models and datasets) in Table 3, we find that LLM-generated UT are better judges for selecting code solutions with higher pass rates as compared to trained, state-of-the-art 8B RM.
UTGen is a Better Judge for Code. Across models and datasets, using UTs from UTGen to select code yields the highest performance, outperforming external RM by 4.36%, randomly generated UTs by 3.73%, and prompted UTs by 1.83% on HE+ with Llama3 8B (with similar trends on MBPP+ and using Qwen2.5 7B).
5.4 On Test-time Scaling and Backtracking
In Sec. 3.3, we highlighted the challenges of debugging with imperfect UTs and suggested remedies to make our debugging pipeline, UTDebug, robust to noisy feedback from such UTs. We study the effectiveness of these measures, i.e., test-time scaling of UT outputs, and backtracking, on debugging with generated UTs for 3 rounds using Qwen2.5 7B on the MBPP+Fix dataset in Table 4.
| UT Method | Acc. | |
| Randomly-sampled (Qwen2.5 7B) | 34.77 | |
| - Output Test-time Scaling | 30.77 | - 4.0 |
| - Backtracking | 32.61 | - 2.2 |
| UTGen (Qwen2.5 7B) | 37.54 | |
| - Output Test-time Scaling | 26.15 | - 11.4 |
| - Backtracking | 34.38 | - 3.2 |
We show that test-time scaling and backtracking are crucial for LLM debugging with generated UTs. From Table 4, we observe that irrespective of the underlying method for UT generation (either randomly sampled or from UTGen), removing either backtracking or test-time scaling hurts downstream performance. First, we find that removing test-time scaling for predicting the output of UTs decreases the performance of randomly-sampled UTs by 4% and that of UTGen by 11.4%. Note that a larger drop in UTGen’s performance when removing test-time scaling of UT output prediction is consistent with the findings in Sec. 5.1 that models struggle with reasoning over correct outputs for failing UT inputs (more often generated by UTGen). Therefore, test-time scaling in output prediction provides a greater boost for reasoning over challenging inputs Wang et al. (2022), and consequently, removing it yields larger drops for UTGen. Moreover, Table 4 demonstrates that, without validation on other generated unit tests, LLMs tend to overfit on the unit test contained in the feedback, resulting in up to 3.2% drop in performance of UTGen without backtracking.
5.5 Scaling with Number of Generated UTs
Thus far, we use generated UTs across baselines and models. However, as we described in Sec. 3.3, having multiple UTs can be advantageous because: (i) there is a higher likelihood of generating a failing UT and getting a reliable signal for when the code is correct, (ii) more robust signal for when to backtrack using validation on the entire generate test suite. We analyze the impact of increasing the number of generated UTs on downstream accuracy of Qwen2.5 7B after 3 rounds of debugging in Fig. 4. Our findings are as follows:
First, Fig. 4 shows that despite increasing the number of generated unit tests, UTGen consistently outperforms randomly sampled UTs that may not be failing. This highlights the benefits of generating targeted unit tests conditioned on buggy code in order to trigger errors and generate appropriate feedback for debugging.
On MBPP+Fix (Hard), which contains less obvious errors and is harder to debug, we find that despite scaling to up to generated UTs, the performance gap between UTGen and randomly-sampled UTs remains at 10% (absolute).
5.6 Comparison to Frontier LLMs
l c c c c Model
Method Attack Rate Out Acc. Acc. Attack GPT-4o Random 23.92 57.25 17.25 Prompted 54.51 60.39 33.33 DeepSeek V3 Random 24.31 63.53 17.84 Prompted 54.31 58.04 33.33 Qwen2.5 32B Random 25.29 47.25 11.37 Prompted 51.96 48.43 29.41 UTGen 56.08 59.22 34.71
In Sec. 5.1 we evaluated open-source LLMs that we trained using UTGen, and in Sec. 3.3 we applied these models in our UTDebug framework for debugging, finding the trained models to be effective. Sec. 5.1 compares stronger frontier models – GPT-4o Achiam et al. (2023) and DeepSeek-V3 Guo et al. (2024b) – to the best open-source model tested (Qwen2.5 32B). Here, we see that even these stronger and much larger frontier models struggle at the task of generating failing unit tests. First, we find that when prompting models with faulty codes, the attack rate of the generated UTs is slightly over 50% on MBPP+Fix (Hard) dataset, showcasing the inherently challenging nature of isolating corner cases and subtle errors in partially correct code. Furthermore, UTGen beats both GPT-4o and DeepSeek-V3 in terms of Acc. Attack, with a gain of over both, as well as in terms of Attack Rate. On output accuracy, UTGen is better than DeepSeek-V3 and slighlty worse than GPT-4o. Overall, we find that using UT-based feedback even with frontier LLMs only triggers and identifies errors 1 in 3 times, leaving tremendous room for growth in the ability of models to identify partially incorrect code.
Note that cost is a major factor in automated debugging; we show that using UTGen for debugging in UTDebug is successful, but involves generating multiple UTs with self-consistency across several rounds of debugging for each problem (cf. Algorithms 1 and 2). Such calls will quickly become costly on frontier models like GPT-4o and DeepSeek-V3, making relatively smaller models trained with UTGen more attractive options.
6 Discussion and Conclusion
Our work on UTGen contributes to the broader landscape of verification and feedback generation for LLM-generated code. While recent work has focused on training verifiers to provide feedback Mahan et al. (2024); Zhang et al. (2024), a key challenge remains in obtaining high-quality feedback signals for debugging. UTGen addresses this by directly generating unit tests that can identify problems in code, complementing existing work on how to effectively incorporate and present feedback for debugging Chen et al. (2023b); Zhong et al. (2024) along with test-time scaling and backtracking incorporated in UTDebug. Our results demonstrate that without a quality signal to determine code correctness and/or how a faulty code is failing (in the form of unit tests), using LLMs to generate feedback and debug still proves to be challenging. This is one of the first efforts in this direction, and we hope to spark more interest in future work toward LLM-generated unit tests (both input and outputs) that reveal the model’s coding errors.
Our approach connects to and complements recent work on handling real-world programming issues. While approaches designed for SWEBench Jimenez et al. (2024) focus on fixing known issues from GitHub by understanding and implementing fixes for bug reports, UTGen addresses the upstream challenge of automatically discovering potential issues in new code through test generation. Both tasks share a core challenge: determining the expected behavior of code without access to correct implementations. This connects to the fundamental concept of simulatability from computability theory Rice (1953), where we ask whether a program can predict the behavior of another program. Recent work such as Jain et al. (2024) shows that while LLMs can often simulate existing code by tracing execution steps, they struggle more with predicting correct outputs from specifications alone. Our results align with these findings – while UTGen can generate test inputs that trigger errors (high attack rate), predicting correct expected outputs remains challenging (lower output accuracy). This suggests that improving LLMs’ ability to reason about intended program behavior from specifications remains a crucial direction for future work. Nevertheless, we find that the modifications made to debugging in UTDebug help boost UTGen’s accuracy and account for noise, leading to downstream gains.
Conclusion.
We first identified a key trade-off between attack rate and output prediction accuracy when predicting unit tests with models. In light of this trade-off, we introduced UTGen, a new method for creating training data and teaching models to produce unit tests. This allows us to train models to produce better unit tests, as measured by intrinsic metrics like attack rate and output accuracy. Moreover, finding that existing data contains large numbers of easy errors, we introduce a new subset of data with challenging and hard-to-diagnose errors. To enable debugging with automated unit tests, we propose UTDebug, wherein we augment our predictor’s accuracy with test-time scaling and regularize it using a cross-validation and back-tracking procedure that prevents it from overfitting to a narrow or incorrect unit test. This, combined with UTGen, results in consistent increases in debugging performance across models, and can be used to improve code generation via best-of-N sampling.
Acknowledgments
We thank Swarnadeep Saha for insightful discussions in the early phases of this work. This work was supported by NSF-CAREER Award 1846185, DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL-2112635, and the Center for AI Safety Compute Cluster. Any opinions, findings, and conclusions or recommendations in this work are those of the author(s) and do not necessarily reflect the views of the sponsors.
Appendix A Experimental Setup Details
A.1 Coding Datasets
We use debugging datasets based on popular LLM coding benchmarks, HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) along with their extended evaluation suites with more unit tests as proposed by Liu et al. (2024b).
HE+Fix. We start with the HumanEvalFix dataset containing human-introduced errors to gold solutions proposed by Muennighoff et al. (2024) (released under MIT license). However, this dataset only uses minimal unit tests to evaluate code correctness (from HumanEval) which has shown to be unreliable Li et al. (2023), as it can miss errors due to low coverage. Therefore, we filter for problems that overlap with EvalPlus Liu et al. (2024b) released under Apache-2.0 license, which contains over 80 more unit tests. This yields 158 problems that each have a faulty solution and an expanded suite of private UTs – i.e., UTs not used for debugging – for evaluation.
MBPP+Fix. We construct a debugging dataset based on the MBPP+ benchmark Austin et al. (2021); Liu et al. (2024b); Ni et al. (2024). To obtain faulty codes, we sample 16 solutions for each problem from different 7-8B scale models and filter for incorrect solutions based on the private unit tests (described in Sec. A.1). Then, we select one faulty solution at random corresponding to each problem. This results in a total of 325 problems with incorrect solutions representing realistic errors made by LLM coders. We find such errors are generally more challenging for models to debug (cf. Sec. 5).